OpenStudy (anonymous):

A line has a slope of and a y-intercept of −14. Which equations represent the line? Choose all answers that are correct.

OpenStudy (anonymous):

a slope of what?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

I think it is a slope of 9/4 and a y-intercept of −14

OpenStudy (anonymous):

a -9 + 4y = -56

OpenStudy (anonymous):

b 9x + 4y = 56

OpenStudy (anonymous):

The slope represents x and y intercept represents y, therefore the slope is y=9/4x -14

OpenStudy (anonymous):

If you put -9x+4y=-56 in a slop intercept form it would be y=9/4x -14